OpenStudy (pythagoras123):

Find the smallest number such that i) it leaves a remainder 2 when divided by 3; ii) it leaves a remainder 3 when divided by 5; iii) it leaves a remainder 5 when divided by 7.

OpenStudy (anonymous):

i)5 ii)8

OpenStudy (anonymous):

iii.)12

OpenStudy (pythagoras123):

The question asks for a number satisfying all the 3 conditions, not 3 separate numbers fufiling each individual condition.

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

LCM of 5,8,12 is the answer

OpenStudy (pythagoras123):

Ermm, LCM of 5,8 and 12 is 120. 120 does not yield any remainder when divided by 3 or 5.

OpenStudy (anonymous):

Sorry for that. I have made a mistake

OpenStudy (hoblos):

68= 3*22 +2 = 5*13 +3 = 7*9 + 5